# The metaplectic semigroup and its applications to time-frequency analysis and evolution operators

### v1

@@ 2.7. Modulation spaces
The estimate (74) follows by choosing
(98) $c=2^{(s-d)/2}c^{\prime}b\int_{{\mathbb{R}^{2d}}}e^{-\pi|z|^{2}}v_{s}(z)dz.$
Finally, by (49), $\widehat{U}^{-1}$ is upper-block triangular if and only if $U$ is upper-block triangular. Using again that $\sigma_{\mathrm{max}}(U^{-1})=\sigma_{\mathrm{max}}(U)$, the estimate (74) reads for $\widehat{U}^{-1}$ as
(99) $\lVert\widehat{U}^{-1}f\rVert_{M^{p,q}_{v_{s}}}=c|\det(D^{\top})|^{1/p-1/q}\sigma_{\mathrm{max}}(U)^{2s}\prod_{j=1}^{d}\Big(\frac{\sigma_{j}}{1+\sigma_{j}^{2}}\Big)^{-1/2}\lVert f\rVert_{M^{p,q}_{v_{s}}},$
where $|\det(D^{\top})|^{1/p-1/q}=1$ if $p=q$, and $|\det(D^{\top})|^{1/p-1/q}=|\det(A)|^{1/q-1/p}$ if $p\neq q$ (for in this case $U^{-1}$ is upper-block triangular). Consequently, inequality (75) follows by applying (99) to $\widehat{U}f$.
@@
(100) $\frac{1}{c}\prod_{j=1}^{d}\Big(\frac{\sigma_{j}}{1+\sigma_{j}^{2}}\Big)^{1/2}\lVert f\rVert_{2}\leq\lVert\widehat{U}f\rVert_{2}\leq c\prod_{j=1}^{d}\Big(\frac{\sigma_{j}}{1+\sigma_{j}^{2}}\Big)^{-1/2}\lVert f\rVert_{2},$
while we know that $\lVert\widehat{S}f\rVert_{2}=\lVert f\rVert_{2}$. A more precise estimate for $p=q=2$ and $s=0$ was obtained in [16, Corollary 3.5] using interpolation results. There, the authors prove
@@
which arises from the term $\langle\widehat{U}g,g\rangle$, and does not depend on $p$ and $q$. We expect sharper estimates can be obtained through Wigner analysis, a finer investigation which is beyond the scope of this manuscript.
@@ 3.1. The positivity condition
It is impossible to extend $\operatorname*{Mp}(d,\mathbb{R})$ to a group of operators, so that $\pi^{Mp}$ extends consistently to a double cover of $\operatorname*{Sp}(d,\mathbb{C})$. A selection of admissible complex symplectic matrices is needed a priori.
@@
where $\sigma$ is the standard complex symplectic form. We denote by $\operatorname*{Sp}_{+}(d,\mathbb{C})$ the semigroup of positive symplectic matrices. Instead, we say that $S$ is strictly positive if (103) holds with >0>0. We refer to (103) as to the Hörmander’s positivity condition.
The semigroup $\operatorname*{Sp}_{+}(d,\mathbb{C})$ is denoted as $\mathcal{C}_{+}$ in [41]. We refer to [41, Lemma 3.2] for a proof of the semigroup property, in the more general framework of positive linear canonical relations.
It will be useful to express Hörmander’s condition (103) more explicitly, as we shall do promptly.
@@
Notice that the condition $S^{\top}JS=J$ defining the symplectic group can be rephrased as
@@
Positive symplectic matrices admit a very beneficial polar-type decomposition, see [41, Proposition 5.10] and [9, Theorem 2.1]. To simplify the next statement and the discussion that follows, let us define the following class of positive symplectic matrices (recall (37) for the definition of $\otimes$).
@@
with $\Xi_{j}=\operatorname*{id}_{L^{2}}$, or $\Xi_{j}=\mathcal{R}_{\vartheta_{j}}$ ($\vartheta_{j}>0$), or $\Xi_{j}=V_{i\alpha_{j}}$ ($\alpha_{j}>0$), $j=1,\ldots,d$.
@@
Although apparently different, [41, Proposition 5.10] and [9, Theorem 2.1] are equivalent. To see it, three points must be discussed.
@@
In Hörmander’s [41, Proposition 5.10], only $\alpha=1$ in (108) is allowed, but this is not restrictive: if $\alpha>0$,
@@
and the matrices $\mathcal{D}_{\alpha^{1/2}},\mathcal{D}_{\alpha^{-1/2}}$ are in $\operatorname*{Sp}(1,\mathbb{R})$, so the contributions of every $\alpha_{j}\neq 1$ can be embodied in $U_{1}$ and $U_{1}$ in (110), respectively.
Finally, in [41, Proposition 5.10], the matrices $\mathcal{R}_{\tau_{j}}$ and $V_{i\alpha_{j}}$ constructing $S_{2}$ can be arranged in any order. However, in [9, Theorem 2.1], the matrix $\Xi$ is constructed by assembling the (conjugate via $\mathcal{W}$ of the blocks) $V_{i\alpha}$ firstly, and the blocks $\mathcal{R}_{\vartheta_{j}}$ secondly. Again, the two statements are equivalent, since the matrices $\mathcal{R}_{\vartheta}\otimes V_{i}$ and $V_{i}\otimes\mathcal{R}_{\vartheta}$ are conjugate through the real symplectic matrix $\mathcal{D}_{L}$, with
@@
[41, Theorem 5.12] A matrix $S\in\operatorname*{Sp}_{+}(d,\mathbb{C})$ is invertible if and only if it is a real symplectic matrix.
For the following result we refer to [9, Theorem 2.5] and [41, Section 5].
@@ 3.2. Definition and main properties of the metaplectic semigroup
see e.g. [45, Theorem 2.6.12]. This operator is bounded on $L^{2}(\mathbb{R}^{d})$. For every $j=1,\ldots,d$, let
(121) $\Xi_{j}=\begin{cases}\mathcal{R}_{\vartheta_{j}}&\text{if $\widehat{S}_{j}=\mathfrak{R}_{\vartheta_{j}}$},\\ V_{i\alpha_{j}}&\text{if $\widehat{S}_{j}=\mathfrak{p}_{i\alpha_{j}}$}.\end{cases}$
@@
The contents of [41, Theorem 5.12] and [9, Proposition 4.2] are summarized in the following Theorem.
@@
(131) $a_{j}=\begin{cases}\cosh\vartheta_{j}&\text{if $S_{j}=\mathcal{R}_{\vartheta_{j}}$},\\ 1&\text{otherwise},\end{cases}\qquad b_{j}=\begin{cases}i\sinh\vartheta_{j}&\text{if $S_{j}=\mathcal{R}_{\vartheta_{j}}$},\\ 0&\text{otherwise},\end{cases}\qquad\lambda_{j}=\begin{cases}\alpha_{j}&\text{if $S_{j}=V_{i\alpha_{j}}$},\\ 0&\text{otherwise}.\end{cases}$
Observe that by construction $\Xi_{1}$ and $\Xi_{2}$ commute, and $\alpha_{j}\vartheta_{j}=0$ for every $j=1,\ldots,d$. We write $\Xi_{1}=\mathcal{R}_{\Theta}$ and $\Xi_{2}=V_{i\Delta}$, and we also denote by $\Theta=(\vartheta_{1},\ldots,\vartheta_{d})$, where we imply $\vartheta_{j}=0$ if $\Xi_{j}$ is the identity or in the form $V_{i\alpha_{j}}$ ($\alpha_{j}>0$), and $\Delta=\operatorname*{diag}(\alpha_{1},\ldots,\alpha_{d})$ where, again, we assume $\alpha_{j}=0$ if $\Xi_{j}=I$ or in the form $\Xi_{j}=\mathcal{R}_{\vartheta_{j}}$ ($\vartheta_{j}>0$). Consequently, if $\widehat{\Xi}$ is an atomic metaplectic contraction, we can write
@@ 4.1. Generators
Now that the metaplectic semigroup has been extensively introduced, we turn to a description of $\operatorname*{Sp}_{+}(d,\mathbb{C})$ and $\operatorname*{Mp}_{+}(d,\mathbb{C})$ that does not resort to atomic metaplectic contractions. At the level of positive symplectic matrices, we shall see that $\operatorname*{Sp}_{+}(d,\mathbb{C})$ is simply obtained by adding matrices $V_{Q}$ with $Q\in\Sigma_{\geq 0}(d)$ to the generators of $\operatorname*{Sp}_{+}(d,\mathbb{R})$.
@@
The semigroup $\operatorname*{Sp}_{+}(d,\mathbb{C})$ is (weakly) generated by $J$ and by the matrices in the form $\mathcal{D}_{E}$, $E\in\operatorname*{GL}(d,\mathbb{R})$, and $V_{Q}$, $Q\in\Sigma_{\geq 0}(d)$.
@@
The matrix $S_{2}\in\mathcal{E}^{\prime}(d)$, so we just need to show that $S_{1}\in\operatorname*{Sp}_{+}^{\prime}(d,\mathbb{C})$. However, by (136),
(137) $S_{1}=\begin{pmatrix}I&O\\ iQ&I\end{pmatrix}\begin{pmatrix}\operatorname*{diag}(a_{1},\ldots,a_{d})&O\\ O&\operatorname*{diag}(a_{1},\ldots,a_{d})^{-1}\end{pmatrix}J^{-1}\begin{pmatrix}I&O\\ iQ&I\end{pmatrix}J,$
with $Q=\operatorname*{diag}(\mu_{1}\ldots,\mu_{d})$, $\mu_{j}=\tanh\vartheta_{j}$ if $S_{j}=\mathcal{R}_{\vartheta_{j}}$, $\mu_{j}=1$ otherwise. This concludes the proof.
@@ 4.2. The polar decomposition
For every linear and bounded operator $T$ on $L^{2}(\mathbb{R}^{d})$ there exist unique operators $T_{1}$ and $T_{2}$ with the following properties [54, Theorem VI.10]:
@@
There exist (non-unique) $\widehat{U}_{1},\widehat{U}_{2}\in\operatorname*{Mp}(d,\mathbb{R})$ and $\widehat{\Xi}$ atomic metaplectic contraction such that $\widehat{S}=\widehat{U}_{1}\widehat{\Xi}\widehat{U}_{2}$. Let $\widehat{U}=\widehat{U}_{1}\widehat{U}_{2}$ and let $\widehat{Z}=\widehat{S}_{2}^{-1}\widehat{\Xi}\widehat{S}_{2}$. By Theorem 3.10 $(vi)$, $\widehat{\Xi}$ is self-adjoint. We may factorize $\widehat{\Xi}=\mathfrak{R}_{\Theta}\mathfrak{p}_{i\Delta}$ as in Remark 3.11. Correspondingly, $\Xi=\mathcal{R}_{\Theta}V_{i\Delta}$ and the two factors also commute. Since $\mathcal{R}_{\vartheta_{j}}^{1/2}=\mathcal{R}_{\vartheta_{j}/2}$ and $V_{i\alpha_{j}}^{1/2}=V_{i\alpha_{j}/2}$, by tensorization we obtain that $\widehat{\Xi}^{1/2}=\mathfrak{R}_{\Theta/2}\mathfrak{p}_{i\Delta/2}$, where we use here that $\mathfrak{R}_{\Theta/2}$ and $\mathfrak{p}_{i\Delta/2}$ commute as well by construction. Since they are also self-adjoint, we infer that for every $f\in L^{2}(\mathbb{R}^{d})$,
@@
Generally speaking, the polar decomposition of $\widehat{S}$ is rather implicit and not always easy to compute. By contrast, atomic metaplectic contractions are explicit. For this reason, resorting to the non-unique factorization in (124) remains useful when proving general statements about metaplectic operators. However, this is no longer sufficient for their phase-space analysis. Indeed, if (144) is the polar decomposition of $\widehat{S}$, then $\widehat{U}$ acts as a sort of real part of $\widehat{S}$, and its action is reflected in a symplectic linear transformation of phase space, as in the classical case. This property fails for $\widehat{Z}$, whose action on phase space must be interpreted differently, see Section 5.2 below.
@@
Thanks to Lemma 4.4 we obtain a characterization of $\operatorname*{Mp}_{0}(d,\mathbb{C})$ and, with it, a full description of $\operatorname*{Mp}_{+}(d,\mathbb{C})$.
@@
Under the notation of Lemma 4.4, let us consider the mapping $\iota:\widehat{S}\in\operatorname*{Mp}_{+}(d,\mathbb{C})\mapsto(\widehat{U},\widehat{Z})\in\operatorname*{Mp}(d,\mathbb{R})\times\operatorname*{Mp}_{0}(d,\mathbb{C})$.
@@
Observe that $\widehat{S}\in\operatorname*{Mp}_{0}(d,\mathbb{C})\cap\operatorname*{Mp}(d,\mathbb{R})$ if and only if $\widehat{S}=c\cdot\operatorname*{id}_{L^{2}}$ for some phase factor $c\in\mathbb{C}$, $|c|=1$. Item $(i)$ is just a trivial restatement of the uniqueness of the polar decomposition, using Lemma 4.4. To prove item $(ii)$, just observe that if $\widehat{Z}\in\operatorname*{Mp}_{0}(d,\mathbb{C})$, then $\widehat{Z}=\widehat{U}_{1}\widehat{\Xi}\widehat{U}_{2}$ as in (124). By basically repeating the argument of Lemma 4.4, we obtain the polar decomposition of $\widehat{Z}$
@@
The other inclusion was already noted in the proof of Lemma 4.4. Item $(iii)$ follows by associativity: set $\widehat{Z}^{\prime}=\widehat{U}_{1}\widehat{\Xi}\widehat{U}_{1}^{-1}$, then

### v2

@@ 2.7. Modulation spaces
This concludes the proof.
@@
(97) $\lVert\widehat{U}f\rVert_{2}\leq c\prod_{j=1}^{d}\Big(\frac{\sigma_{j}}{1+\sigma_{j}^{2}}\Big)^{-1/2}\lVert f\rVert_{2},$
while we know that $\lVert\widehat{S}f\rVert_{2}=\lVert f\rVert_{2}$. A more precise estimate for $p=q=2$ and $s=0$ was obtained in [15, Corollary 3.5] using interpolation results. There, the authors prove
@@
with exponent 1/2-1/2, which arises from the term $\langle\widehat{U}g,g\rangle$, and does not depend on $p$ and $q$. We expect sharper estimates can be obtained through Wigner analysis, a finer investigation which is beyond the scope of this manuscript.
Remark 2.14.
Another formulation of (74) can be obtained by using [19, Proposition 105] in (93). This gives
(100) $|\langle\widehat{U}g,g\rangle|=|\det(I_{d}-i\mathcal{C}(U)^{-1})|^{-1/2},$
where $\mathcal{C}(U)=(C+iD)(A+iB)^{-1}$ is the Cayley transform of $U$, see also [27]. Consequently,
(101) $\lVert\widehat{U}f\rVert_{M^{p,q}_{v_{s}}}\lesssim|\det(A)|^{1/p-1/q}\sigma_{\mathrm{max}}(U)^{2s}|\det(I_{d}-i\mathcal{C}(U)^{-1})|^{1/2}\lVert f\rVert_{M^{p,q}_{v_{s}}}.$
@@ 3.1. The positivity condition
It is impossible to extend $\operatorname*{Mp}(d,\mathbb{R})$ to a group of operators that are bounded on $L^{2}(\mathbb{R}^{d})$, so that $\pi^{Mp}$ extends consistently to a double cover of $\operatorname*{Sp}(d,\mathbb{C})$. A selection of admissible complex symplectic matrices is needed a priori.
@@
where $\sigma$ is the complexified standard symplectic form. We denote by $\operatorname*{Sp}_{+}(d,\mathbb{C})$ the semigroup of positive symplectic matrices. Instead, we say that $S$ is strictly positive if (102) holds with >0>0.
The semigroup $\operatorname*{Sp}_{+}(d,\mathbb{C})$ is denoted as $\mathcal{C}_{+}$ in [39]. We refer to [39, Lemma 3.2] for a proof of the semigroup property, in the more general framework of positive linear canonical relations.
It will be useful to express condition (102) more explicitly, as we shall do promptly.
@@
Notice that the condition $S^{\top}JS=J$ defining $\operatorname*{Sp}(d,\mathbb{C})$ can be rephrased in terms of $\mathcal{M}_{S}$ as
@@
Positive symplectic matrices admit a very beneficial polar-type decomposition, see [39, Proposition 5.10] and [9, Theorem 2.1]. To simplify the next statement and the discussion that follows, let us define the following class of positive symplectic matrices (recall (37) for the definition of $\otimes$).
@@
with $\Xi_{j}=I_{2\times 2}$, or $\Xi_{j}=\mathcal{R}_{\vartheta_{j}}$ ($\vartheta_{j}>0$), or $\Xi_{j}=V_{i\alpha_{j}}$ ($\alpha_{j}>0$), $j=1,\ldots,d$.
@@
Although apparently different, [39, Proposition 5.10] and [9, Theorem 2.1] are equivalent. To see it, three points must be discussed.
@@
In Hörmander’s [39, Proposition 5.10], only $\alpha=1$ in (107) is allowed, but this is not restrictive: if $\alpha>0$,
@@
and the matrices $\mathcal{D}_{\alpha^{1/2}},\mathcal{D}_{\alpha^{-1/2}}$ are in $\operatorname*{Sp}(1,\mathbb{R})$, so the contributions of every $\alpha_{j}\neq 1$ can be embodied in $U_{1}$ and $U_{2}$ in (109), respectively.
Finally, in [39, Proposition 5.10], the matrices $\mathcal{R}_{\tau_{j}}$ and $V_{i\alpha_{j}}$ constructing $\Xi$ can be arranged in any order. However, in [9, Theorem 2.1], the matrix $\Xi$ is constructed by assembling the (conjugate via $\mathcal{W}$ of the blocks) $V_{i\alpha}$ firstly, and the blocks $\mathcal{R}_{\vartheta_{j}}$ secondly. Again, the two constructions are equivalent, since the matrices $\mathcal{R}_{\vartheta}\otimes V_{i}$ and $V_{i}\otimes\mathcal{R}_{\vartheta}$ are conjugate through the real symplectic matrix $\mathcal{D}_{L}$, with
@@
[39, Theorem 5.12] A matrix $S\in\operatorname*{Sp}_{+}(d,\mathbb{C})$ is invertible if and only if it is a real symplectic matrix.
For the following result we refer to [9, Theorem 2.5] and [39, Section 5].
@@ 3.2. Definition and main properties of the metaplectic semigroup
see e.g. [43, Theorem 2.6.12]. This operator is bounded on $L^{2}(\mathbb{R}^{d})$. For every $j=1,\ldots,d$, let
(120) $\Xi_{j}=\begin{cases}\mathcal{R}_{\vartheta_{j}}&\text{if $\widehat{\Xi}_{j}=\mathfrak{R}_{\vartheta_{j}}$},\\ V_{i\alpha_{j}}&\text{if $\widehat{\Xi}_{j}=\mathfrak{p}_{i\alpha_{j}}$}.\end{cases}$
@@
The contents of [39, Theorem 5.12] and [9, Proposition 4.2] are summarized in the following Theorem.
@@
(130) $a_{j}=\begin{cases}\cosh\vartheta_{j}&\text{if $\Xi_{j}=\mathcal{R}_{\vartheta_{j}}$},\\ 1&\text{otherwise},\end{cases}\qquad b_{j}=\begin{cases}i\sinh\vartheta_{j}&\text{if $\Xi_{j}=\mathcal{R}_{\vartheta_{j}}$},\\ 0&\text{otherwise},\end{cases}\qquad\lambda_{j}=\begin{cases}\alpha_{j}&\text{if $\Xi_{j}=V_{i\alpha_{j}}$},\\ 0&\text{otherwise}.\end{cases}$
Observe that by construction $\Xi_{1}$ and $\Xi_{2}$ commute, and $\alpha_{j}\vartheta_{j}=0$ for every $j=1,\ldots,d$. We write $\Xi_{1}=\mathcal{R}_{\Theta}$ and $\Xi_{2}=V_{i\Delta}$, with $\Theta=(\vartheta_{1},\ldots,\vartheta_{d})$, where we imply $\vartheta_{j}=0$ if $\Xi_{j}$ is the identity or in the form $V_{i\alpha_{j}}$ ($\alpha_{j}>0$), and $\Delta=\operatorname*{diag}(\alpha_{1},\ldots,\alpha_{d})$ where, again, we assume $\alpha_{j}=0$ if $\Xi_{j}=I$ or in the form $\Xi_{j}=\mathcal{R}_{\vartheta_{j}}$ ($\vartheta_{j}>0$). Consequently, if $\widehat{\Xi}$ is an atomic metaplectic contraction, we can write
@@ 4.1. Generators
Now that the metaplectic semigroup has been extensively introduced, we turn to a description of $\operatorname*{Sp}_{+}(d,\mathbb{C})$ and $\operatorname*{Mp}_{+}(d,\mathbb{C})$ that does not resort to atomic metaplectic contractions. At the level of positive symplectic matrices, we shall see that $\operatorname*{Sp}_{+}(d,\mathbb{C})$ is simply obtained by adding matrices $V_{Q}$ with $Q\in\Sigma_{\geq 0}(d)$ to the generators of $\operatorname*{Sp}(d,\mathbb{R})$.
@@
The semigroup $\operatorname*{Sp}_{+}(d,\mathbb{C})$ is generated by $J$ and by the matrices in the form $\mathcal{D}_{E}$, $E\in\operatorname*{GL}(d,\mathbb{R})$, and $V_{Q}$, $Q\in\Sigma_{\geq 0}(d)$.
@@
The matrix $V_{i\Delta}\in\mathcal{E}^{\prime}(d)$, so we just need to show that $\mathcal{R}_{\Theta}\in\operatorname*{Sp}_{+}^{\prime}(d,\mathbb{C})$. By (135),
(136) $\mathcal{R}_{\Theta}=\begin{pmatrix}I&O\\ iQ&I\end{pmatrix}\begin{pmatrix}\operatorname*{diag}(a_{1},\ldots,a_{d})&O\\ O&\operatorname*{diag}(a_{1},\ldots,a_{d})^{-1}\end{pmatrix}J^{-1}\begin{pmatrix}I&O\\ iQ&I\end{pmatrix}J,$
with $Q=\operatorname*{diag}(\mu_{1}\ldots,\mu_{d})$, $\mu_{j}=\tanh\vartheta_{j}$ if $\Xi_{j}=\mathcal{R}_{\vartheta_{j}}$, $\mu_{j}=1$ otherwise. This concludes the proof.
@@ 4.2. The polar decomposition
For every linear and bounded operator $T$ on $L^{2}(\mathbb{R}^{d})$ there exist unique operators $T_{1}$ and $T_{2}$ with the following properties [52, Theorem VI.10]:
@@
There exist (non-unique) $\widehat{U}_{1},\widehat{U}_{2}\in\operatorname*{Mp}(d,\mathbb{R})$ and $\widehat{\Xi}$ atomic metaplectic contraction such that $\widehat{S}=\widehat{U}_{1}\widehat{\Xi}\widehat{U}_{2}$. Let $\widehat{U}=\widehat{U}_{1}\widehat{U}_{2}$ and let $\widehat{Z}=\widehat{U}_{2}^{-1}\widehat{\Xi}\widehat{U}_{2}$. By Theorem 3.10 $(vi)$, $\widehat{\Xi}$ is self-adjoint. We may factorize $\widehat{\Xi}=\mathfrak{R}_{\Theta}\mathfrak{p}_{i\Delta}$ as in Remark 3.11. Correspondingly, $\Xi=\mathcal{R}_{\Theta}V_{i\Delta}$ and the two factors also commute. Since $\mathcal{R}_{\vartheta_{j}}^{1/2}=\mathcal{R}_{\vartheta_{j}/2}$ and $V_{i\alpha_{j}}^{1/2}=V_{i\alpha_{j}/2}$, by tensorization we obtain that $\widehat{\Xi}^{1/2}=\mathfrak{R}_{\Theta/2}\mathfrak{p}_{i\Delta/2}$, where we use here that $\mathfrak{R}_{\Theta/2}$ and $\mathfrak{p}_{i\Delta/2}$ commute as well by construction. Since they are also self-adjoint, we infer that for every $f\in L^{2}(\mathbb{R}^{d})$,
@@
Generally speaking, the polar decomposition of $\widehat{S}$ is rather implicit and not always easy to compute. Specifically,
(145) $\widehat{Z}=(\widehat{S}^{\ast}\widehat{S})^{1/2},\qquad\widehat{U}=\widehat{S}(\widehat{S}^{\ast}\widehat{S})^{-1/2}.$
In contrast, atomic metaplectic contractions are explicit. For this reason, resorting to the non-unique factorization in (123) remains useful when proving general statements about metaplectic operators. However, this is no longer sufficient for their phase-space analysis. Indeed, if (143) is the polar decomposition of $\widehat{S}$, then $\widehat{U}$ acts as a sort of real part of $\widehat{S}$, and its action is reflected in a symplectic linear transformation of phase space, as in the classical case. This property fails for $\widehat{Z}$, whose action on phase space must be interpreted differently, see Section 5.2 below.
@@
Thanks to Theorem 4.4 we obtain a characterization of $\operatorname*{Mp}_{0}(d,\mathbb{C})$ and, with it, a full description of $\operatorname*{Mp}_{+}(d,\mathbb{C})$.
@@
Under the notation of Theorem 4.4, let us consider the mapping $\iota:\widehat{S}\in\operatorname*{Mp}_{+}(d,\mathbb{C})\mapsto(\widehat{U},\widehat{Z})\in\operatorname*{Mp}(d,\mathbb{R})\times\operatorname*{Mp}_{0}(d,\mathbb{C})$.
@@
Observe that $\widehat{S}\in\operatorname*{Mp}_{0}(d,\mathbb{C})\cap\operatorname*{Mp}(d,\mathbb{R})$ if and only if $\widehat{S}=c\cdot\operatorname*{id}_{L^{2}}$ for some phase factor $c\in\mathbb{C}$, $|c|=1$. Item $(i)$ is just a trivial restatement of the uniqueness of the polar decomposition, using Theorem 4.4. To prove item $(ii)$, just observe that if $\widehat{Z}\in\operatorname*{Mp}_{0}(d,\mathbb{C})$, then $\widehat{Z}=\widehat{U}_{1}\widehat{\Xi}\widehat{U}_{2}$ as in (123). By basically repeating the argument of Theorem 4.4, we obtain the polar decomposition of $\widehat{Z}$
@@
The other side of the equivalence was already noted in the proof of Theorem 4.4. Item $(iii)$ follows by associativity: set $\widehat{Z}^{\prime}=\widehat{U}_{1}\widehat{\Xi}\widehat{U}_{1}^{-1}$, then